OpenStudy (anonymous):

The sum of the reciprocals of two consecutive odd integers is 11 over 60. Find the integers.

OpenStudy (whpalmer4):

I think you have not stated the problem correctly...

OpenStudy (anonymous):

oh yeah... it should be:The sum of the reciprocals of two consecutive even integers is \[\frac{ 11 }{ 60 }\]. Find the integers.

OpenStudy (whpalmer4):

Okay, let the first integer be \(x\), and the second integer will be \(x+2\). \[\frac{1}{x}+\frac{1}{x+2} =\frac{11}{60}\]sharpen your pencil and do the algebra :-)

OpenStudy (whpalmer4):

you should actually get two solutions, but only one of them will make sense for this problem